# “Fear will keep the local systems in line.”

### v2

@@ 0.1. What was done in previous papers?
Namely, this work proves that
@@ 0.4. What is not done in this paper?
  • Geometric Langlands for Fargues-Fontaine curves.
@@ 1.1. The Langlands functor and its categorical properties
In [GLC1, Sect. 1], we have constructed a functor
@@ 1.2. Spectral properties
  • The functor ${\mathbb{L}}_{G}$ is ${\operatorname{QCoh}}(\operatorname{LS}_{\check{G}})$-linear, i.e., it intertwines the actions of ${\operatorname{QCoh}}(\operatorname{LS}_{\check{G}})$ on the two sides, see [GLC1, Sect. 1.7]. As ${\operatorname{QCoh}}(\operatorname{LS}_{\check{G}})$ is rigid monoidal, this implies that the functor ${\mathbb{L}}_{G}^{L}$ is also ${\operatorname{QCoh}}(\operatorname{LS}_{\check{G}})$-linear;
@@ 2. Reduction to the case when $G$ is almost simple and simply-connected
For a number of technical reasons, [^2] One reason is not very serious: if $G$ has a center of positive dimension, the stack $\operatorname{LS}^{\operatorname{irred}}_{\check{G}}$ is not smooth, which is a silly annoyance. The more serious reason is that when $g=2$ we will need to assume that $G$ has no factors of $A_{1}$ in its Dynkin diagram, see the preamble to Sect. 5.5. in the main argument in the proof of GLC, we will need to assume that the group $G$ is almost simple and simply-connected. In this section we will perform a reduction to this case.
@@ 2.1. Compatibility between Langlands functors
Let $G_{1}$ and $G_{2}$ be a pair of reductive groups, and let
@@
given by restriction along $\phi^{\vee}$.
@@
First, we can factor $\phi$ as
$G_{1}\overset{\phi^{\prime}}{\to}G_{1}\times Z^{0}_{G_{2}}\overset{\phi^{\prime\prime}}{\to}G_{2}.$
The proposition evidently holds for $\phi^{\prime}$. Hence, it is sufficient to prove it for $\phi^{\prime\prime}$. Thus, we can assume that our $\phi$ is surjective.
2.2.2.
When $\phi$ is surjective, we factor it as
$G_{1}\overset{\phi^{\prime}}{\to}G_{1}/\operatorname{ker}(\phi)^{0}\overset{\phi^{\prime\prime}}{\to}G_{2},$
so that the kernel of $\phi^{\prime}$ is a connected torus, and $\phi^{\prime\prime}$ is an actual isogeny.
The case of a surjection with kernel being a connected torus will be analyzed explicitly in Sect. 2.4.2.
Hence, it is enough to prove Proposition 2.1.6 for $\phi$ being an isogeny (this assumption will be used in one place in the course of the proof, namely, in Sect. 2.2.4).
2.2.3.
@@
where the horizontal composite functors are
@@
By [GLC1, Theorem 1.6.2], the functors ${\mathbb{L}}_{G_{i},\operatorname{coarse}}$ send
@@
The statement for the clockwise circuit follows similarly: indeed since $\phi:G_{1}\to G_{2}$ is an isogeny, the map
@@
is étale, and hence the functor
$\phi^{!}:\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}})\to\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}})$
preserves compactness.
@@
Once we prove that the outer circuit in (2.3) commutes as a diagram of ${\operatorname{QCoh}}(\operatorname{LS}_{{\check{G}}_{1}})$-module categories, it would follow that its restriction to compact objects commutes as a diagram of $\operatorname{Perf}(\operatorname{LS}_{{\check{G}}_{1}})$-module categories, so the resulting left commuting square in (2.3) is one of $\operatorname{Perf}(\operatorname{LS}_{{\check{G}}_{1}})$-module categories, and hence of $\operatorname{Ind}(\operatorname{Perf}(\operatorname{LS}_{{\check{G}}_{1}}))\simeq{\operatorname{QCoh}}(\operatorname{LS}_{{\check{G}}_{1}})$-module categories.
@@
is fully faithful, it suffices to show that the outer square in the concatenation of (2.4) with the commutative diagram
@@
is the functor of co-induction, i.e., the right adjoint to the functor
@@
Note that pullback with respect to the map
@@
induces a functor
$\phi^{!}:\operatorname{Whit}(G_{2})\to\operatorname{Whit}(G_{2}).$
It is easy to see that the following diagram commutes:
@@
Combining this with the commutative diagrams [GLC2, Equation (18.4)]
@@
for $i=1,2$, we obtain that the required commutativity is equivalent to the commutativity of
@@
2.2.9.
@@ 2.3. Changing the group
Indeed, this is particular case (extended to stacks in a straightforward way) of [AG, Corollary 7.6.2]. In Sect. 2.4, we will prove:
@@ 2.4.3. Proof for type A
and it is easy to see that the functor $(\phi^{!})^{\prime}$ is fully faithful. In fact, its essential image is a direct summand in $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{1}})$, described as follows:
@@
Now,
@@
where $\alpha$ runs over the subset consisting of those characters that vanish on the subgroup
$\operatorname{ker}(Z_{G_{1}}\to Z_{G_{2}}/Z^{0}_{G_{2}}).$
The functor $\phi_{!}$ factors as
where the first arrow is the corresponding orthogonal projection.
2.4.4.
@@
Let us assume this proposition and proceed with the proof of Case A of Theorem 2.3.4. We obtain that the functor (2.8) is an equivalence if and only if the functor
@@
2.4.6.
Consider ${\operatorname{QCoh}}(\operatorname{LS}_{\check{T}})$ as a monoidal category with respect to convolution. As such, it acts on ${\operatorname{QCoh}}(\operatorname{LS}_{{\check{G}}_{2}})$ via the map
$\psi^{\vee}:\operatorname{LS}_{\check{T}}\to\operatorname{LS}_{{\check{G}}_{2}}.$
@@
2.4.7.
Note the action of ${\operatorname{QCoh}}(\operatorname{LS}_{{\check{G}}_{2}})$ on $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}})$ is compatible with the actions of ${\operatorname{QCoh}}(\operatorname{LS}_{\check{T}})$ on both. Thus, (2.14) can be rewritten as the special case (for ${\mathbf{C}}:=\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}})$) of the functor
@@
2.4.8.
We claim that (2.15) is an equivalence for any such $\mathbf{C}$}. Indeed, let $\mathcal{Y}$} be an algebraic stack with an affine diagonal, and let
@@
Assume that both $\mathcal{Y}$} and $\mathcal{T}$} are quasi-compact, locally almost of finite type and eventually coconnective (so that [Ga3, Theorem 2.2.6] is applicable).

### v3

@@ 0.1. What was done in previous papers?
Namely, that work proves that
@@ 0.4. What is not done in this paper?
  • Geometric Langlands for the Fargues-Fontaine curve.
@@ 1.1. The Langlands functor and its categorical properties
In [GLC1, Sect. 1], we constructed a functor
@@ 1.2. Spectral properties
  • The functor ${\mathbb{L}}_{G}$ is ${\operatorname{QCoh}}(\operatorname{LS}_{\check{G}})$-linear, i.e., it intertwines the actions of ${\operatorname{QCoh}}(\operatorname{LS}_{\check{G}})$ on the two sides, see [GLC1, Sect. 1.7]. As the monoidal category ${\operatorname{QCoh}}(\operatorname{LS}_{\check{G}})$ is rigid, this implies that the functor ${\mathbb{L}}_{G}^{L}$ is also ${\operatorname{QCoh}}(\operatorname{LS}_{\check{G}})$-linear;
@@ 2. Reduction to the case when $G$ is almost simple and simply-connected
For a number of technical reasons [^2] One reason is not very serious: if $G$ has a center of positive dimension, the stack $\operatorname{LS}^{\operatorname{irred}}_{\check{G}}$ is not smooth, which is a silly annoyance. The more serious reason is that when $g=2$ we will need to assume that $G$ has no factors of $A_{1}$ in its Dynkin diagram, see the preamble to Sect. 5.5., in the main argument in the proof of GLC, we will need to assume that the group $G$ is almost simple (by which we mean simple modulo a finite center) and simply-connected. In this section we will perform a reduction to this case.
@@ 2.1. Compatibility between Langlands functors
Let $G_{1}$ and $G_{2}$ be a pair of reductive groups (over our ground field $k$), and let
@@
given by restriction along $\phi^{\vee}$, see Sect. 2.2.11.
@@
where the composite horizontal functors are
@@
2.2.3.
By the construction in [GLC1, Corollary 1.6.5], the functors ${\mathbb{L}}_{G_{i}}$ send
@@
We now prove the statement for the clockwise circuit. Note that the map
@@
is not schematic, so the functor
(2.4) $\phi^{!}:\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}})\to\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}}).$
does not preserve compactness. To circumvent this, we argue as follows.
First, we note that (2.4) does preserve compactness if the initial homomorphism $\phi$ has a finite kernel. So in this case, the clockwise circuit does send $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}})^{c}$ to ${\operatorname{IndCoh}}_{\operatorname{Nilp}}(\operatorname{LS}_{{\check{G}}_{1}})^{>-\infty}$.
Thus, let us temporarily assume having proved Proposition 2.1.6 for such homomorphisms.
2.2.4.
In general, we decompose $\phi$ as
$G_{1}\overset{\phi^{\prime}}{\to}G^{\prime}_{2}\overset{\psi}{\to}G_{2},$
where
$G^{\prime}_{2}:=G_{2}\times G_{1,\operatorname{ab}},$
and $\psi$ is the projection. The assertion of Proposition 2.1.6 trivially holds for $\psi$.
Note now that the homomorphism $\phi^{\prime}$ does have a finite kernel. Hence, by assumption, the functor ${\mathbb{L}}_{G_{1}}\circ(\phi^{\prime})^{!}$ is isomorphic to $((\phi^{\prime})^{\vee})_{*}\circ{\mathbb{L}}_{G^{\prime}_{2}}$, and hence
${\mathbb{L}}_{G_{1}}\circ\phi^{!}\simeq{\mathbb{L}}_{G_{1}}\circ(\phi^{\prime})^{!}\circ\psi^{!}\simeq((\phi^{\prime})^{\vee})_{*}\circ{\mathbb{L}}_{G^{\prime}_{2}}\circ\psi^{!}\simeq((\phi^{\prime})^{\vee})_{*}\circ(\psi^{\vee})_{*}\circ{\mathbb{L}}_{G_{2}}\simeq(\phi^{\vee})_{*}\circ{\mathbb{L}}_{G_{2}}.$
In particular, the clockwise circuit for $\phi$ also sends
$\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}})^{c}\to{\operatorname{IndCoh}}_{\operatorname{Nilp}}(\operatorname{LS}_{{\check{G}}_{1}})^{>-\infty}.$
@@
Once we prove that the outer circuit in (2.3) commutes as a diagram of ${\operatorname{QCoh}}(\operatorname{LS}_{{\check{G}}_{1}})$-module categories, it would follow that its restriction to compact objects commutes as a diagram of $\operatorname{Perf}(\operatorname{LS}_{{\check{G}}_{1}})$-module categories, so the resulting left commuting square in (2.3) is one of $\operatorname{Perf}(\operatorname{LS}_{{\check{G}}_{1}})$-module categories, and hence of $\operatorname{Ind}(\operatorname{Perf}(\operatorname{LS}_{{\check{G}}_{1}}))\simeq{\operatorname{QCoh}}(\operatorname{LS}_{{\check{G}}_{1}})$-module categories [^3] The latter equivalence follows from the fact that $\operatorname{LS}_{\check{G}}$ admits a (global) embedding into a smooth stack, see [AG, Sect. 10.6.2 and Proposition 10.6.6(a)]..
@@
is fully faithful (see [GLC4, Proposition 1.1.4]), it suffices to show that the outer square in the concatenation of (2.5) with the commutative diagram
@@
is the functor of co-induction [^4] It may be more common to call this operation just “induction”., i.e., the right adjoint to the functor
@@
The map (of factorization ind-schemes)
@@
induces a pair of adjoint functors
(2.8) $\phi_{!}:\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{1}})\rightleftarrows\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{2}}):\phi^{!}$
(as factorization categories), which in turn induce adjoint functors
$\phi_{!}:\operatorname{Whit}(G_{1})\rightleftarrows\operatorname{Whit}(G_{2}):\phi^{!}$
(also as factorization categories).
2.2.9.
In addition, we claim that the functors (2.8) also induce (factorization) functors
$\operatorname{Sph}_{G_{1}}\simeq\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{1}})^{{\mathfrak{L}}^{+}(G_{1})}\overset{{}^{\prime}\!\phi^{!}}{\leftarrow}\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{2}})^{{\mathfrak{L}}^{+}(G_{2})}\simeq\operatorname{Sph}_{G_{2}}$
and
$\operatorname{Sph}_{G_{1}}\simeq\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{1}})^{{\mathfrak{L}}^{+}(G_{1})}\overset{{}^{\prime}\!\phi_{!}}{\to}\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{2}})^{{\mathfrak{L}}^{+}(G_{2})}\simeq\operatorname{Sph}_{G_{2}}$
compatible with the forgetful functors $\operatorname{Sph}_{G_{i}}\to\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{i}})$.
Namely, the functor $^{\prime}\!\phi^{!}$ is
$\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{2}})^{{\mathfrak{L}}^{+}(G_{2})}\to\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{2}})^{{\mathfrak{L}}^{+}(G_{1})}\overset{\phi^{!}}{\to}\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{1}})^{{\mathfrak{L}}^{+}(G_{1})},$
where the first arrow is the functor of forgetting ${\mathfrak{L}}^{+}(G_{2})$-equivariance to ${\mathfrak{L}}^{+}(G_{1})$-equivariance.
The functor $^{\prime}\!\phi_{!}$ is
$\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{1}})^{{\mathfrak{L}}^{+}(G_{1})}\overset{\phi_{!}}{\to}\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{2}})^{{\mathfrak{L}}^{+}(G_{1})}\to\operatorname{D-mod}_{\frac{1}{2}}({\operatorname{Gr}}_{G_{2}})^{{\mathfrak{L}}^{+}(G_{2})},$
where the second arrow is pullback with respect to the projection
@@
equal to
@@
where:
  • The 2nd isomorphism is obtained from the identification ${\mathfrak{L}}^{+}(G_{2})\simeq{\mathfrak{L}}^{+}(Z_{G_{2}})\overset{{\mathfrak{L}}^{+}(Z_{G_{1}})}{\times}{\mathfrak{L}}^{+}(G_{1})$;
  • The 3rd isomorphism is obtained from noticing that the action of $\operatorname{pt}/{\mathfrak{L}}^{+}(Z_{G_{1}})$ on the double quotient ${\mathfrak{L}}^{+}(G_{1})\backslash{\mathfrak{L}}(G_{2})/{\mathfrak{L}}^{+}(G_{2})$ on the left equals the action obtained from the projection $\operatorname{pt}/{\mathfrak{L}}^{+}(Z_{G_{1}})\to\operatorname{pt}/{\mathfrak{L}}^{+}(Z_{G_{2}})$, and the action of $\operatorname{pt}/{\mathfrak{L}}^{+}(Z_{G_{2}})$ on ${\mathfrak{L}}^{+}(G_{1})\backslash{\mathfrak{L}}(G_{2})/{\mathfrak{L}}^{+}(G_{2})$ on the right;
  • The 4th arrow is induced by the multiplication map
    $\operatorname{pt}/{\mathfrak{L}}^{+}(Z_{G_{2}})\overset{\operatorname{pt}/{\mathfrak{L}}^{+}(Z_{G_{1}})}{\times}\operatorname{pt}/{\mathfrak{L}}^{+}(Z_{G_{2}})\to\operatorname{pt}/{\mathfrak{L}}^{+}(Z_{G_{2}}).$
Remark 2.2.10.
Note that the functor $^{\prime}\!\phi_{!}$ is not the left adjoint of $^{\prime}\!\phi^{!}$: the latter would involve an additional step of !-averaging from ${\mathfrak{L}}^{+}(G_{1})$-equivariance to ${\mathfrak{L}}^{+}(G_{2})$-equivariance.
Note also that the functor $^{\prime}\!\phi_{!}$ is not monoidal.
2.2.11.
It follows from the construction of the naive geometric Satake functor $\operatorname{Sat}_{G}^{\operatorname{nv}}$ (see [Ras, Sect. 6.28-6.35]) that the diagram (of factorization functors)
@@
commutes.
Unwinding, we obtain that in the global situation, the functor
$\phi^{!}:\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}})\to\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{1}})$
is compatible with the actions of ${\mathop{\operatorname{\rm Rep}}}({\check{G}}_{1})_{\operatorname{Ran}}$, where the action on $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}})$ is via $\operatorname{Res}^{\phi^{\vee}}$.
2.2.12.
By the definition of the Casselman-Shalika equivalence $\operatorname{CS}_{G}$ (see [GLC2, Sect. 1.4]), the commutativity of (2.9) implies that the diagram (of factorization functors)
[…]
also commutes.
Since the horizontal arrows in the latter diagram are equivalences, we obtain that the diagram obtained by passing to right adjoints along the vertical arrows also commutes:
[…]
2.2.13.
Combining this with [^5] We recall that the subscript “Ran” in the next few formulas indicates that we are taking global sections of the corresponding crystal of categories over the Ran space, see [GLC2, Sect. B.11.1]. the commutative diagrams [GLC2, Equation (18.4)]
[…]
for $i=1,2$, we obtain that the commutativity of (2.7) is equivalent to the commutativity of
[…]
2.2.14.
@@ 2.3. Changing the group
Indeed, this is particular case of [AG, Corollary 7.6.2] (extended to stacks in a straightforward way [^6] One uses the fact that for any affine scheme $S$ mapping to $\operatorname{LS}_{{\check{G}}_{i}}$, the operation ${\operatorname{QCoh}}(S)\underset{{\operatorname{QCoh}}(\operatorname{LS}_{{\check{G}}_{i}})}{\otimes}$- commutes with limits on ${\operatorname{QCoh}}(\operatorname{LS}_{{\check{G}}_{i}})$-module categories.).
2.3.4.
In Sect. 2.4 below, we will prove:
@@ 2.4.3. Proof for type A
and it is easy to see that the functor $(\phi^{!})^{\prime}$ is fully faithful. In fact, its essential image is a direct summand in $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{1}})$, described as follows.
2.4.4.
@@
We can identify $G_{1}$ with the derived group of $G_{2}$; moreover
$G_{2}\simeq G_{1}\overset{\Gamma}{\times}Z^{0}_{G_{2}},$
where $\Gamma$ is a finite group equipped with embeddings
(2.17) $Z_{G_{1}}\hookleftarrow\Gamma\hookrightarrow Z^{0}_{G_{2}}.$
We claim:
Lemma 2.4.5.
@@
where $\alpha$ runs over the subset consisting of those characters that vanish on $\Gamma$.
Proof.
The subcategory $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{1}})^{\prime}$ is the full subcategory in $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{1}})$, generated by the essential image of $\phi^{!}$.
Denote
$\operatorname{Bun}^{\prime}_{G_{1}}:=\operatorname{Bun}_{G_{1}}\overset{\operatorname{pt}/\Gamma}{\times}(\operatorname{pt}/Z^{0}_{G_{2}}),$
so that we can factor $\phi$ as
$\operatorname{Bun}_{G_{1}}\overset{\phi^{\prime}}{\to}\operatorname{Bun}_{G^{\prime}_{1}}\to\operatorname{Bun}_{G_{2}},$
where the second arrow is a closed embedding. Hence, it suffices to show that the essential image of $(\phi^{\prime})^{!}$ generates the subcategory described in the statement of the lemma.
Consider the projection
$\operatorname{Bun}^{\prime}_{G_{1}}\overset{\phi^{\prime\prime}}{\to}\operatorname{Bun}_{G_{1}}/(\operatorname{pt}/\Gamma).$
It is a gerbe with typical fiber $\operatorname{pt}/Z^{0}_{G_{2}}$. Hence, the essential image of $(\phi^{\prime\prime})^{!}$ generates $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}^{\prime}_{G_{1}})$.
Hence, the subcategory generated by the essential image of $(\phi^{\prime})^{!}$ is the same as the one generated by the essential image of $(\phi^{\prime\prime}\circ\phi^{\prime})^{!}$. However, the latter is exactly the category singled out by the condition on the characters, since $\phi^{\prime\prime}\circ\phi^{\prime}$ is the projection
$\operatorname{Bun}_{G_{1}}\to\operatorname{Bun}_{G_{1}}/(\operatorname{pt}/\Gamma).$
2.4.6.
@@
Note that we have a commutative diagram
[…]
see Sect. 5.1.3, and note that
$G_{2}/Z^{0}_{G_{2}}\simeq G_{1}/\Gamma\,\Rightarrow\,Z_{G_{2}/Z^{0}_{G_{2}}}\simeq Z_{G_{1}}/\Gamma,$
where $\Gamma$ is as in (2.17).
@@
2.4.8.
Let us assume this proposition and proceed with the proof of Case A of Theorem 2.3.5. Note that the functor $\phi_{!}$ factors as
where the first arrow is the corresponding orthogonal projection.
We obtain that the functor (2.12) is an equivalence if and only if the functor
@@
2.4.9.
Consider ${\operatorname{QCoh}}(\operatorname{LS}_{\check{T}})$ as a monoidal category with respect to convolution. As such, it acts on ${\operatorname{QCoh}}(\operatorname{LS}_{{\check{G}}_{2}})$ by convolution, corresponding to the action of $\operatorname{LS}_{\check{T}}$ on $\operatorname{LS}_{{\check{G}}_{2}}$ given by the map
$\psi^{\vee}:{\check{T}}\to Z_{{\check{G}}_{2}}.$
@@
2.4.10.
Note the action of ${\operatorname{QCoh}}(\operatorname{LS}_{{\check{G}}_{2}})$ on $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}})$ is compatible with the actions of ${\operatorname{QCoh}}(\operatorname{LS}_{\check{T}})$ on both: indeed, this is a particular case of the compatibility in Sect. 2.1.3 for the homomorphism $G_{2}\to G_{2}\times T$.
Thus, (2.19) can be rewritten as the special case (for ${\mathbf{C}}:=\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G_{2}})$) of the functor
@@
2.4.11.
We claim that (2.20) is an equivalence for any such $\mathbf{C}$}. Here is the general paradigm:
Let $\mathcal{Y}$} be an algebraic stack with an affine diagonal, and let
@@
Assume that both $\mathcal{Y}$} and $\mathcal{T}$} are quasi-compact, locally almost of finite type and eventually coconnective (so that [Ga3, Theorem 2.2.6] [^7] Some gaps in the proof of this theorem have been found after its publication, which have subsequently been fixed. However, for our purposes here, we are only using the case of algebraic stacks that can be written as global quotients, in which case the result follows from the (easier) Theorem 2.2.4 in loc.cit. is applicable).
@@
2.4.12.
We apply this to
${\mathcal{Y}}:=\operatorname{LS}^{\prime}_{{\check{G}}_{1}},\,\,\widetilde{\mathcal{Y}}:=\operatorname{LS}_{{\check{G}}_{2}},\,\,{\mathcal{T}}:=\operatorname{LS}_{{\check{T}}}.$